If each edge of a cube is doubled, 
how many times will its volume increase?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the edge of cube be“l”.
Volume of cube, V =l^3
When edge of cube is doubled, then

Volume of cube, say V’= (2l)^3= 8( l^3)
V’ = 8×V
Hence, volume will increase 8 times.

Thanks